# Sequence-aware Pre-training for Echocardiography Probe Movement Guidance

### v2

@@ Abstract
Echocardiography is an essential medical technique for diagnosing cardiovascular diseases, but its high operational complexity has led to a shortage of trained professionals. To address this issue, we introduce a novel probe movement guidance algorithm that has the potential to be applied in guiding robotic systems or novices with probe pose adjustment for high-quality standard plane image acquisition. Cardiac ultrasound faces two major challenges: (1) the inherently complex structure of the heart, and (2) significant individual variations. Previous works have only learned the population-averaged structure of the heart rather than personalized cardiac structures, leading to a performance bottleneck. Clinically, we observe that sonographers dynamically adjust their interpretation of a patient’s cardiac anatomy based on prior scanning sequences, consequently refining their scanning strategies. Inspired by this, we propose a novel sequence-aware self-supervised pre-training method. Specifically, our approach learns personalized three-dimensional cardiac structural features by predicting the masked-out image features and probe movement actions in a scanning sequence. We hypothesize that if the model can predict the missing content it has acquired a good understanding of personalized cardiac structure. Extensive experiments on a large-scale expert scanning dataset with 1.31 million samples demonstrate that our proposed sequence-aware paradigm can effectively reduce probe guidance errors compared to other advanced baseline methods. Our code will be released after acceptance.
@@ 1 Introduction
Cardiovascular disease is the leading threat to human health [1]. The most common method for assessing cardiac health is transthoracic echocardiography, as recommended by the American Society of Echocardiography [2]. In this procedure, a sonographer manually maneuvers a probe to find critical two-dimensional planes and observes the myocardium, valves, and blood flow [3, 4, 5] in these planes (views) in real-time using ultrasound images. However, due to the complexity of the heart’s structure, even small changes in the 6 degrees of freedom (6-DOF) pose of the probe can affect the imaging of the structures being observed. Furthermore, individual differences in heart size, shape, and structure [6, 7] make it even more challenging, requiring adjustments to scanning strategies based on each patient’s unique characteristics. This places high demands on the sonographer’s scanning skills, leading to a long learning curve for sonographers and a shortage of practitioners [8].
@@
Previous researchers proposed assistive scanning systems [9, 10], which provide 6-DOF pose adjustment suggestions for the probe based on the current ultrasound image to reach the target plane. These system has the potential to improve the quality and efficiency with which inexperienced sonographers find standard planes, thereby increasing the global supply of echocardiography. These works have attempted to model the population-averaged cardiac structure from large-scale data and rely solely on a single plane’s images to make action decisions. This paradigm neglects individual variations in cardiac structure, leading to a performance bottleneck. Clinically, as shown in Fig. 1, sonographers learn personalized cardiac structural information based on the acquired planes and their spatial pose information from the patient, further assisting in subsequent scans. From this observation, we can derive two key insights: (1) sonographers make decisions based on previous sequence data (plane images and corresponding probe pose) rather than a single image, and (2) sequence data contains personalized information about cardiac structure. Several researchers [11, 12] have also explored using sequences as network inputs. However, we empirically find that simply increasing the sequence length does not lead to continuous performance improvement (Sec. 4). Therefore, models does not inherently possess the ability to extract beneficial individual cardiac structural information from sequences. To this end, we propose a novel sequence-aware self-supervised pre-training method, which learns personalized cardiac structures through a masked prediction task. To be specific, we sample a scanning trajectory comprising plane images and the probe movement actions between them, which are derived from the probe’s pose, from the same patient. Both the plane images and actions are processed by modality-specific encoders to obtain corresponding features, with some features randomly masked as targets for prediction. For the masked images, we apply supervision in the feature space, as [13] demonstrated that this is a more efficient learning method. The actions, on the other hand, are re-predicted to their original values for supervision. This paradigm requires the model to predict the post-action plane visual feature based on the given probe movement action and visible patient’s cardiac structure context information, or to predict the probe movement action between two planes. Thus, this method enables the model to acquire the ability to construct an implicit three-dimensional cardiac structure based on known patient information.
@@
  • We introduce a cardiac probe movement guidance model that aims to assist in acquiring high-quality standard plane images. To the best of our knowledge, this is the first guidance model that is technically disclosed and capable of covering up to ten critical standard planes.
@@
  • Extensive experiments conducted on a large-scale dataset containing 1.31 million clinical expert-scanned samples demonstrate that our method consistently outperforms baseline approaches across various experimental settings, providing more accurate guidance signals.
@@ 2 Related Work
With the advancement of AI technologies [14, 15, 16], AI-assisted applications have demonstrated significant potential in echocardiography [9, 10, 17, 18].
@@
AI-assisted ultrasound scanning. To achieve AI-assisted scanning, one line of research adopts reinforcement learning methods [19, 20], while another focuses on the supervised learning paradigm [9, 10, 11, 12]. For reinforcement learning methods, the primary challenge lies in constructing a good simulation environment. Current approaches [19, 20] mainly use computed tomography (CT) scan data of the heart as the simulation environment. However, acquiring cardiac CT scan data is more costly than ultrasound data, which limits the scalability of such methods. Additionally, differences between CT and ultrasound data pose challenges for domain adaptation. For supervised learning methods [9, 11, 10, 12], expert scanning demonstration data is typically collected, where each plane image is paired with a corresponding probe pose. This data is used to learn the mapping between plane images and the corresponding probe movement actions. For instance, Narang et al. [9] proposed a single-frame based commercial AI algorithm, trained on 500K labeled samples, to guide novices in performing echocardiography. However, this is a clinical-oriented work, and the technical details have not been fully disclosed. Unlike them, we propose a novel algorithm and make their details publicly available. Bao et al. [10] proposed a navigation system that provides only probe rotation movement guidance for locating the apical four-chamber (A4C) and two-chamber (A2C) plane. However, considering only the A4C and A2C planes is far from sufficient to meet clinical needs. Transitioning from the A4C plane to the parasternal long-axis plane involves a complex combination of translational and rotational motions, making it highly restrictive to focus solely on rotation. Furthermore, Bao et al. did not consider the issue of individual differences. Meanwhile, our paper proposes a novel pre-training method to enhance the model’s ability to construct personalized cardiac structures, while also providing transition and rotation guidance for locating the ten most common standard planes, including the A4C and A2C planes. Droste et al. [11] proposed US-GuideNet, which uses sequence as input to guide obstetric ultrasound scanning. Empirically, we find that simply increasing the sequence length does not lead to performance improvements, indicating that the model’s ability to capture individual structural information from sequences needs optimization (Sec. 4). However, US-GuideNet does not focus on enhancing the model’s ability to extract useful information from past sequences, resulting in suboptimal performance. Moreover, US-GuideNet also focuses exclusively on rotation, which represents a major limitation of its work.
@@
AI-assisted ultrasound diagnosis. Unlike scanning, AI-assisted ultrasound diagnosis [21, 22, 23, 24, 25] has developed much faster due to the abundance of available data. For instance, Christensen et al. [21] proposed a vision-language foundation model EchoCLIP for echocardiogram interpretation. EchoCLIP is a strong cardiac diagnosis model trained with over 1 million video-text pairs from 224K echocardiography studies across a decade of clinical care. Meanwhile, Jiao et al. proposed USFM [22], a universal ultrasound foundation model designed for multi-task and multi-organ applications. These powerful diagnostic foundation models learn visual features useful for ultrasound scanning, so we include them in comparisons.
@@ 3 Dataset and method
Figure 3: Echocardiographic scanning dataset. (a) Dataset statistics. (b) Dataset collection system. (c) Collected ten standard planes. Images are sourced from the guideline [2].
@@ 3.1 Echocardiographic Scanning Dataset
To conduct the research, we collected a large-scale dataset comprising 356 expert scanning trajectories from 178 adult subjects, performed by two sonographers with over 10 years of experience. To collect the data, we constructed a system as shown in Fig. 3b, where sonographers manipulated a probe attached to the end of a robotic arm to perform continuous scans of ten crucial echocardiogram planes (Fig. 3c, recommended by the American Society of Echocardiography [2]). The system simultaneously recorded the images captured by the probe and the corresponding 6 degrees of freedom (6-DOF) pose data, forming a scan sequence $\{(\mathbf{I}_{t},\mathbf{p}_{t})\}^{T}_{t=1}$, where $t$ is the timestamp, $\mathbf{I}_{t}\in\mathbb{R}^{C\times H\times W}$ is the ultrasound image at time $t$, and $\mathbf{p}_{t}\in\mathbb{R}^{6}$ is the corresponding probe pose. The probe’s 6-DOF pose is represented by the translational coordinates along the $x$, $y$, and $z$ axes, and the corresponding rotational Euler angles around the $x$, $y$, and $z$ axes. The values of translation (mm) and rotation (°) are of the same order of magnitude. Since the probe belongs to the rigid transformation group, we can calculate the relative movement action $\mathbf{a}_{i\rightarrow j}$ between two probe poses $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$.
@@
3.2 Sequence-aware Pre-training
@@
Input. Firstly, given an image $I_{t}$ from a specific scan, we sample $L-1$ additional images and their corresponding poses from the same scan. The sampling strategy will be explained in Sec. 3.3. The $L-1$ sampled images are sorted in ascending order based on their timestamps, with $I_{t}$ positioned at the end of the sequence as $I_{t_{L}}$. Then, their relative movement actions are calculated based on their poses, thus generating a scanning trajectory:
@@
$\mathbf{f}_{t_{i}}^{\mathrm{v}}$ $=\mathbf{T}_{i}+\frac{1}{N}\sum_{n=1}^{N}\mathbf{H}_{t_{i},n}^{\mathrm{v}},$ $\mathbf{f}_{t_{i}}^{\mathrm{v}},\mathbf{T}_{i}\in\mathbb{R}^{D},$ (3)
@@
$\mathbf{f}_{t_{j}}^{\mathrm{a}}$ $=\mathbf{T}_{j}+A_{\phi}(\mathbf{a}_{t_{j}\rightarrow t_{j+1}}),$ $\mathbf{f}_{t_{j}}^{\mathrm{a}},\mathbf{T}_{j}\in\mathbb{R}^{D},$ (4)
@@
$\mathbf{m}_{t_{q}}^{\mathrm{v}^{{}^{\prime}}}$ $=\mathbf{m}_{t_{q}}^{\mathrm{v}}+\mathbf{T}_{q}+\frac{1}{N}\sum_{n=1}^{N}\mathbf{E}_{n}^{\mathrm{pos}},$ $\mathbf{m}_{t_{q}}^{\mathrm{v}}\in\mathbb{R}^{D},$ (5)
$\mathbf{m}_{t_{k}}^{\mathrm{a}^{{}^{\prime}}}$ $=\mathbf{m}_{t_{k}}^{\mathrm{a}}+\mathbf{T}_{k},$ $\mathbf{m}_{t_{k}}^{\mathrm{a}}\in\mathbb{R}^{D},$ (6)
@@ 3.3 Sampling Strategy
Bidirectional protocol. In clinical practice, novice physicians typically begin by scanning relatively simple planes before transitioning to more challenging ones. When handling complex planes, it is essential to gather as much information as possible about the patient’s cardiac structure to assist in the scanning process. Thus, we also consider a bidirectional sampling method, which captures the maximum amount of plane information (Fig. 5(a2)).
Segmental sampling. The previous two methods define the sampling intervals, within which we apply the segmental sampling strategy [26], as shown in Fig. 5(b). This method first divides the sampling interval into $L-1$ equal segments and then randomly samples one point from each segment with equal probability. This method ensures sampling coverage across the entire interval, enhancing the diversity of sampled planes to capture as much cardiac structural information as possible.
@@ 3.4 Downstream Transfer
[…]
@@
During fine-tuning, the historical sequence sampling protocol remains consistent with pre-training. In unidirectional protocol, when predicting the movement from $I_{t_{L}}$ to standard planes with timestamp $t>t_{L}$, the historical sequence is sampled from $t\leq t_{L}$. Similarly, for standard planes with timestamp $t<t_{L}$, the sequence is segmentally sampled from $t\geq t_{L}$. In bidirectional protocol, during training, the sampling process is the same as pre-training. During evaluation, when predicting the movement from $I_{t_{L}}$ to a specific standard plane, both the labeled standard plane and samples with similar visual features which are identified using a standard plane classifier are removed from the scan. By filtering out these samples, information leakage during evaluation can be avoided. The historical sequence is then segmentally sampled from the remaining instances inside the scan. Importantly, during evaluation, for both unidirectional and bidirectional protocols, the step-2 of segmental sampling (Fig. 5) is replaced with a fixed strategy that selects the sample at the center of each segment, thereby avoiding the introduction of randomness into the validation process.

### v3

# UltraSeP: Sequence-aware Pre-training for Echocardiography Probe Movement Guidance

@@ Abstract
Echocardiography is an essential medical technique for diagnosing cardiovascular diseases, but its high operational complexity has led to a shortage of trained professionals. To address this issue, we introduce UltraSeP, a novel probe movement guidance algorithm that has the potential to be applied in guiding robotic systems or novices with probe pose adjustment for high-quality standard plane image acquisition. Cardiac ultrasound faces two major challenges: (1) the inherently complex structure of the heart, and (2) significant individual variations. Previous works have only learned the population-averaged structure of the heart rather than personalized cardiac structures, leading to a performance bottleneck. Clinically, we observe that sonographers dynamically adjust their interpretation of a patient’s cardiac anatomy based on prior scanning sequences, consequently refining their scanning strategies. Inspired by this, we propose a novel sequence-aware self-supervised pre-training method. Specifically, our approach learns personalized three-dimensional cardiac structural features by predicting the masked-out image features and probe movement actions in a scanning sequence. We hypothesize that if the model can predict the missing content it has acquired a good understanding of personalized cardiac structure. Extensive experiments on a large-scale expert scanning dataset with 1.67 million samples demonstrate that our proposed sequence-aware paradigm can effectively reduce probe guidance errors compared to other advanced baseline methods.
@@ 1 Introduction
Cardiovascular disease is the leading threat to human health [1]. The most common method for assessing cardiac health is transthoracic echocardiography, as recommended by the American Society of Echocardiography [2]. In this procedure, a sonographer manually maneuvers a probe to find critical two-dimensional planes and observes the myocardium, valves, and blood flow [3] in these planes (views) in real-time using ultrasound images. However, due to the complexity of the heart’s structure, even small changes in the 6 degrees of freedom (6-DOF) pose of the probe can affect the imaging of the structures being observed. Furthermore, individual differences in heart size, shape, and structure [4] make it even more challenging, requiring adjustments to scanning strategies based on each patient’s unique characteristics. This places high demands on the sonographer’s scanning skills, leading to a long learning curve for sonographers and a shortage of practitioners [5].
@@
Previous researchers proposed assistive scanning systems [6, 7, 8], which provide 6-DOF pose adjustment suggestions for the probe based on the current ultrasound image to reach the target plane. These system has the potential to improve the quality and efficiency with which inexperienced sonographers find standard planes, thereby increasing the global supply of echocardiography. These works have attempted to model the population-averaged cardiac structure from large-scale data and rely solely on a single plane’s images to make action decisions. This paradigm neglects individual variations in cardiac structure, leading to a performance bottleneck. Clinically, as shown in Fig. 1, sonographers learn personalized cardiac structural information based on the acquired planes and their spatial pose information from the patient [9, 10], further assisting in subsequent scans. From this observation, we can derive two key insights: (1) sonographers make decisions based on previous sequence data (plane images and corresponding probe pose) rather than a single image, and (2) sequence data contains personalized information about cardiac structure. Several researchers [11, 12] have also explored using sequences as network inputs. However, we empirically find that simply increasing the sequence length does not lead to continuous performance improvement (Sec. 4). Therefore, models does not inherently possess the ability to extract beneficial individual cardiac structural information from sequences. To this end, we propose a novel sequence-aware self-supervised pre-training method, namely UltraSeP, which learns personalized cardiac structures through a masked prediction task. To be specific, we sample a scanning trajectory comprising plane images and the probe movement actions between them, which are derived from the probe’s pose, from the same patient. Both the plane images and actions are processed by modality-specific encoders to obtain corresponding features, with some features randomly masked as targets for prediction. For the masked images, we apply supervision in the feature space, as [13] demonstrated that this is a more efficient learning method. The actions, on the other hand, are re-predicted to their original values for supervision. This paradigm requires the model to predict the post-action plane visual feature based on the given probe movement action and visible patient’s cardiac structure context information, or to predict the probe movement action between two planes. Thus, this method enables the model to acquire the ability to construct an implicit three-dimensional cardiac structure based on known patient information.
@@
  • We introduce a cardiac probe movement guidance model (UltraSeP) that aims to assist in acquiring high-quality standard plane images. To the best of our knowledge, this is the first guidance model that is technically disclosed and capable of covering up to ten critical standard planes.
@@
  • Extensive experiments conducted on a large-scale dataset containing 1.67 million clinical expert-scanned samples demonstrate that our method consistently outperforms baseline approaches across various experimental settings, providing more accurate guidance signals.
@@ 2 Related Work
With the advancement of AI technologies [14, 15, 16], AI-assisted applications have shown significant potential in echocardiography, such as ejection fraction estimation [17], and anatomical structure segmentation [18, 19]
AI-assisted ultrasound scanning. To achieve AI-assisted scanning, one line of research adopts reinforcement learning methods [20, 21, 22], while another focuses on the supervised learning paradigm [7, 11, 23].
For reinforcement learning methods, the primary challenge lies in constructing a good simulation environment. Current approaches [20, 21, 22] mainly use computed tomography (CT) scan data of the heart as the simulation environment. For example, Amadou et al. [20] first segments anatomical structures from the CT data to identify the cardiac region, and then reconstructs this region into an ultrasound volume as training environment. A similar strategy is adopted in works [21]. However, acquiring cardiac CT scan data is more costly than ultrasound data, which limits the scalability of such methods. Additionally, differences between CT and ultrasound data pose challenges for domain adaptation. Bi et al. [22] also attempted to artificially construct a simulation environment for vascular organs. However, such manually constructed simulation environments struggle to model the anatomical variations present in real clinical settings, particularly for organs like the heart, which exhibit significant inter-individual variability. Therefore, although reinforcement learning strategies hold significant potential, their development in ultrasound-assisted scanning remains relatively limited due to the difficulty in constructing realistic simulation environments.
@@
For supervised learning methods, expert scanning demonstration data is typically collected, where each plane image is paired with a corresponding probe pose or adjustment action. This data is used to learn the mapping between plane images and the corresponding probe movement actions. Recent works [23] has demonstrated that this strategy shows promising generalization performance in vascular scanning. Specifically, Narang et al. [7] proposed a single-frame based commercial AI algorithm, trained on 500K labeled samples, to guide novices in performing echocardiography. However, this is a clinical-oriented work, and the technical details have not been fully disclosed. Unlike them, we propose a novel algorithm and make their details publicly available. Bao et al. [8] proposed a navigation system that provides only probe rotation movement guidance for locating the apical four-chamber (A4C) and two-chamber (A2C) plane. However, considering only the A4C and A2C planes is far from sufficient to meet clinical needs. Besides, transitioning from the A4C plane to the parasternal long-axis plane involves a complex combination of translational and rotational motions, making it highly restrictive to focus solely on rotation. Jiang et al. [6, 24] proposed pre-training strategies to learn the implicit cardiac structure, which is helpful for the downstream probe guidance task. Furthermore, Jiang et al. did not consider the issue of individual differences. Meanwhile, our paper proposes a novel pre-training method to enhance the model’s ability to construct personalized cardiac structures. Droste et al. [11] proposed US-GuideNet and Men et al. [12] proposed Multimodal-GuideNet. Both approaches use sequence as input to guide obstetric ultrasound scanning. Empirically, we find that simply increasing the sequence length does not lead to performance improvements, indicating that the model’s ability to capture individual structural information from sequences needs optimization (Sec. 4). However, US-GuideNet and Multimodal-GuideNet do not focus on enhancing the model’s ability to extract useful information from past sequences, resulting in suboptimal performance. Moreover, they also focuses exclusively on rotation, which represents a major limitation of these works.
@@
AI-assisted ultrasound diagnosis. Unlike scanning, AI-assisted ultrasound diagnosis [25, 26, 27, 28, 29] has developed much faster due to the abundance of available data. For instance, Christensen et al. [25] proposed a vision-language foundation model EchoCLIP for echocardiogram interpretation. EchoCLIP is a strong cardiac diagnosis model trained with over 1 million video-text pairs from 224K echocardiography studies across a decade of clinical care. Meanwhile, Jiao et al. proposed USFM [26], a universal ultrasound foundation model designed for multi-task and multi-organ applications. These two powerful diagnostic foundation models [25, 26] learn visual features useful for ultrasound scanning, so we include them in the experimental section for comparison.
@@ 3 Dataset and method
Figure 3: Echocardiographic scanning dataset. (a) Dataset statistics. The dataset was collected by two senior sonographers who performed scans on 238 adult subjects, two scans per individual. This resulted in a large-scale dataset containing 1.67 million samples. (b) Dataset collection system. The sonographer operated a probe attached to the end of a robotic arm for scanning. The system records the ultrasound images along with the corresponding probe poses. (c) Collected ten standard planes. The sonographer scanned six standard planes from the parasternal window and four standard planes from the apical window. Images are sourced from the guideline [2].
@@ 3.1 Echocardiographic Scanning Dataset
To conduct the research, we collected a large-scale dataset comprising 476 expert scanning trajectories from 238 adult subjects, performed by two sonographers with over 10 years of experience. To collect the data, we constructed a system as shown in Fig. 3b, where sonographers manipulated a probe attached to the end of a robotic arm to perform continuous scans of ten crucial echocardiogram planes (Fig. 3c, recommended by the American Society of Echocardiography [2]). The system simultaneously recorded the images captured by the probe and the corresponding 6 degrees of freedom (6-DOF) pose data, forming a scan sequence $\{(\mathbf{I}_{t},\mathbf{p}_{t})\}^{T}_{t=1}$, where $t$ is the timestamp, $\mathbf{I}_{t}\in\mathbb{R}^{C\times H\times W}$ is the ultrasound image at time $t$, and $\mathbf{p}_{t}\in\mathbb{R}^{6}$ is the corresponding probe pose. The probe’s 6-DOF pose is represented by the translational coordinates along the $x$, $y$, and $z$ axes, and the corresponding rotational Euler angles around the $x$, $y$, and $z$ axes. The values of translation (mm) and rotation (°) are of the same order of magnitude. Since the probe belongs to the rigid transformation group, we can calculate the relative movement action $\mathbf{a}_{i\rightarrow j}$ between two probe poses $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$.
@@
3.2 UltraSeP: Sequence-aware Pre-training
@@
Input. Firstly, given an image $I_{t}$ from a specific scan, we sample L1L-1 additional images and their corresponding poses from the same scan. The sampling strategy will be explained in Sec. 3.3. The L1L-1 sampled images are sorted in ascending order based on their timestamps, with $I_{t}$ positioned at the end of the sequence as $I_{t_{L}}$. Then, their relative movement actions are calculated based on their poses, thus generating a scanning trajectory:
@@
$\mathbf{f}_{t_{i}}^{\mathrm{v}}$ $=\mathbf{T}_{i}+\frac{1}{N}\sum_{n=1}^{N}\mathbf{H}_{t_{i},n}^{\mathrm{v}},\ \$$\mathbf{f}_{t_{i}}^{\mathrm{v}},\mathbf{T}_{i}\in\mathbb{R}^{D},$ (3)
@@
$\mathbf{f}_{t_{j}}^{\mathrm{a}}$ $=\mathbf{T}_{j}+A_{\phi}(\mathbf{a}_{t_{j}\rightarrow t_{j+1}}),\ \$$\mathbf{f}_{t_{j}}^{\mathrm{a}},\mathbf{T}_{j}\in\mathbb{R}^{D},$ (4)
@@
$\mathbf{m}_{t_{q}}^{\mathrm{v}^{{}^{\prime}}}$ $=\mathbf{m}_{t_{q}}^{\mathrm{v}}+\mathbf{T}_{q}+\frac{1}{N}\sum_{n=1}^{N}\mathbf{E}_{n}^{\mathrm{pos}},\ \$$\mathbf{m}_{t_{q}}^{\mathrm{v}}\in\mathbb{R}^{D},$ (5)
$\mathbf{m}_{t_{k}}^{\mathrm{a}^{{}^{\prime}}}$ $=\mathbf{m}_{t_{k}}^{\mathrm{a}}+\mathbf{T}_{k},\ \$$\mathbf{m}_{t_{k}}^{\mathrm{a}}\in\mathbb{R}^{D},$ (6)
@@ 3.3 Sampling Strategy
Bidirectional protocol. In clinical practice, novice physicians typically begin by scanning relatively simple planes before transitioning to more challenging ones. When handling complex planes, it is essential to gather as much information as possible about the patient’s cardiac structure to assist in the scanning process. Thus, we also consider a bidirectional sampling method, which captures the maximum amount of plane information (Fig. 5 a2).
Segmental sampling. The previous two methods define the sampling intervals, within which we apply the segmental sampling strategy [30], as shown in Fig. 5(b). This method first divides the sampling interval into L1L-1 equal segments and then randomly samples one point from each segment with equal probability. This method ensures sampling coverage across the entire interval, enhancing the diversity of sampled planes to capture as much cardiac structural information as possible.
@@ 3.4 Downstream Transfer
After pre-training, the sequence-aware transformer $S_{\psi}$ is further fine-tuned for the probe movement guidance task. The goal of the probe movement guidance task is to predict the relative movement action from any arbitrary plane to the ten standard planes. During fine-tuning, we introduce ten guidance prediction heads $G_{\xi_{g}},g\in\mathbb{Z},\ 1\leq g\leq 10$ to predict the action toward each standard plane, as illustrated in Fig. 6. All ten heads are trained simultaneously during fine-tuning, while during inference, the corresponding head is selected based on the target standard plane. Meanwhile, the sequence-aware transformer is shared among the ten heads, as the cardiac structural information it learns is applicable for navigating toward all standard planes. Similar to pre-training, given an image $I_{t_{L}}$ from a specific scan, we sample L1L-1 instances to form a historical sequence. Formally, the forward process for predicting action from $I_{t_{L}}$ to $g$-th standard plane can be represented as:
@@
During fine-tuning, the historical sequence sampling protocol remains consistent with pre-training. In unidirectional protocol, when predicting the movement from $I_{t_{L}}$ to standard planes with timestamp $t>t_{L}$, the historical sequence is sampled from $t\leq t_{L}$. Similarly, for standard planes with timestamp $t<t_{L}$, the sequence is segmentally sampled from $t\geq t_{L}$. In bidirectional protocol, during training, the sampling process is the same as pre-training. During evaluation, when predicting the movement from $I_{t_{L}}$ to a specific standard plane, both the labeled standard plane and samples with similar visual features which are identified using a standard plane classifier are removed from the scan. By filtering out these samples, information leakage during evaluation can be avoided. Otherwise, if the target plane is present in the sequence, its known pose would allow direct derivation of the action, undermining the evaluation of the model’s true capability. The historical sequence is then segmentally sampled from the remaining instances inside the scan. Importantly, during evaluation, for both unidirectional and bidirectional protocols, the step-2 of segmental sampling (Fig. 5) is replaced with a fixed strategy that selects the sample at the center of each segment, thereby avoiding the introduction of randomness into the validation process.